Is it right... To say.... dy/dx = dx/dy

Grade:12

7 Answers

Swapnil Saxena
12 years ago

No this is incorrect.

(dy/dx) is actually equal to (dx/dy)-1

Chetan Mandayam Nayakar
12 years ago

dy/dx=1/(dx/dy)

Menka Malguri
12 years ago

No,dy/dx=1/(dx/dy)

kunal ramteke
12 years ago

i dont know

Rohan Das
10 years ago

acctualy dy/dx=1/dx/dy

 

Akash Kumar Dutta
10 years ago

no

Abhishekh kumar sharma
10 years ago

never

dy\dx is change of y w.r.t x

but dx\dy is change of x w.r.t y which will turn out to be inverse of each other